OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (1, 1). @Sav_1012 @Savannah_Noelle

OpenStudy (anonymous):

But the options are: (2, -5) (-2, 5) (2, 5) (-2, -5)

OpenStudy (savannah_noelle):

y=2x+1

OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (4, -5).

OpenStudy (savannah_noelle):

-5=-3(4) -5=-12+b 7+b

OpenStudy (savannah_noelle):

*7=b